### v3

@@ 6 Future Direction
Large language models (LLMs) are increasingly influential across various industries, and the rise of privatized and domain-specific LLMs presents new topics for management research. This work represents the inaugural attempt to train open-source large language models to the domain of optimization modeling, liberating the field from reliance on proprietary advanced models, which facilitates the application and commercialization of large models within industrial sectors. We hope this work shed lights on the deep integration of large language models with operations research and management science, potential future research directions are listed as follows:
  • Incorporating Sophisticated Techniques. Building upon our foundational work in addressing the scarcity of training data, future research could delve into more advanced techniques. These might include training multiple specialized large language models (LLMs) for collaborative multi-agent modeling, refining transformer architectures to enhance reasoning abilities, and utilizing alignment algorithms like reinforcement learning from human feedback (RLHF), direct preference optimization (DPO), and Kahneman-Tversky optimization (KTO) to boost the performance of large models. Such investigations could unlock the potential of large models to surpass traditional methods in tackling complex operations research issues. Additionally, these strategies could be adapted to address emerging industrial challenges, such as predictive tasks in machine learning.
  • Knowledge Management and Human-Machine Collaboration. Large language models (LLMs) exhibit superior learning efficiency and expansive knowledge storage, heralding a new era in human-computer collaboration and enterprise management. LLMs can leverage technologies such as retrieval-augmented generation (RAG) for historical knowledge management, effectively preserving corporate memory and mitigating the impact of employee turnover. Their advanced text mining capabilities have the potential to revolutionize market research analysis and report generation. This leads to human-machine interactions that transcend the knowledge encapsulated within trained LLMs by integrating a company’s accumulated historical experiences and external human-derived information. Such integration facilitates more efficient task execution, such as optimization modeling, thereby driving digital transformation and fostering innovation within enterprises.
  • Consumer Behavior Insights through LLM. Large models’ capability to simulate human social behaviors offers valuable insights for enterprises analyzing consumer behavior and formulating market strategies. For instance, through simulations, large language models can generate preferences that closely resemble human behaviors, assisting businesses in conducting comprehensive A/B testing. Additionally, the inherent choice models in algorithms such as RLHF facilitate the capture of consumer choice behaviors, providing new perspectives on revenue management issues, including product assortment.
  • Economic and Game Theoretical Aspects of LLM. The commercialization of large models introduces new research dimensions in management studies. The emergence of large models transforms specific industry textual data, such as project reports, into a valuable new data asset. Managing, pricing, and transacting these data assets remain unresolved challenges. Additionally, the advent of large models creates new game relationships between R&D companies and service companies, employees and LLMs, as well as LLMs and customers, related problems in this field like contract theory, mechanism design, and game theory are still open.
@@ 7 Conclusion
This research addresses the issue of insufficient training data in optimization modeling. Looking ahead, we plan to explore potential enhancements through diverse architectures, agent-based models, and advanced modeling techniques to elevate modeling capabilities, leveraging the training dataset provided by OR-Instruct. Additionally, we will examine managerial innovations and practical challenges arising from the real-world application of large language models in industrial contexts.
Acknowledgement
This research is partially supported by National Natural Science Foundation of China (Grants 72394360, 72394364, 72171141), the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing. Zizhuo Wang is partially supported by the National Natural Science Foundation of China (NSFC) [Grant NSFC-72150002, NSFC-72394361] and the Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001). Dongdong Ge is partially supported by the National Natural Science Foundation of China (NSFC) [Grant NSFC-72150001, 72225009, 72394360, 72394365].

### v4

@@ 5.2 Ablation Study
Ablation Study on Question Types
ORLM can be tailored to specific domains within the field of optimization, depending on its initial training dataset. A key consideration in this process is whether incorporating data from other domains is necessary. This investigation aims to provide valuable guidance for implementing ORLM in specialized areas.
We use linear programming as a case study, comparing ORLM’s performance when trained solely on linear programming data versus a mixed dataset including other problem types. Specifically, we construct two training sets, each comprising 7,049 instances: one set exclusively consisting of pure linear programming (LP) problems, and the other comprising a diverse mixture of problem typesThis dataset is comprised of 35% linear programming, 40% mixed integer programming, 25% other types.. We fine-tune the Meta-Llama3 8B model on both datasets, ensuring that hyper-parameter settings remain consistent across experiments. The test performance is summarized in Table 6:
Table 6: Performance of ORLM-LLaMA-3-8B on different datasets
Dataset Data Size NL4OPT MAMO EasyLP MAMO ComplexLP IndustryOR Micro Avg Macro Avg
LP Only 7,049 82.0% 77.4% 26.1% 20.0% 65.4% 51.4%
Mixed 7,049 86.5% 81.2% 34.6% 21.0% 69.8% 55.8%
As illustrated in Table 6, even with identical test sizes, the inclusion of mixed problem types significantly enhances the performance of the ORLM-LLaMA-3-8B across various datasets. Notably, when examining the tests focused on linear programming, specifically MAMO-EasyLP and MAMO-ComplexLP, it is evident that integrating additional problem types is more effective in enhancing the ORLM’s performance on challenging linear programming problems. A plausible explanation for this phenomenon is that the incorporation of a broader variety of problems effectively strengthens the generalization ability of the ORLM-LLaMA-3-8B, allowing it to capture a wider modeling capacity and consequently improve modeling performance (See dong2023abilities).
5.3 Scaling Laws
The concept of scaling laws is a key element in the development of large models. It refers to the phenomenon that the performance of large language models improves as their size and complexity increase. In this section, we will examine the impact of scaling laws on ORLMs from two dimensions: training data and model size.
From the perspective of training data, we begin with seed data and subsequently randomly sample a predetermined number of training data from the data generated by OR-Instruct. Using the LLaMA-3-8B model as the base, we conduct training while keeping all other hyperparameters constant (as in Table 1), in order to observe the performance of ORLM-LLaMA-3-8B on various test sets. Regarding model size, we focus on the Qwen large model series, as it offers a more refined range of model sizes that are compatible with our hardware limitations (0.5B, 1.5B, 3B, 7B, 14B). We fix the training dataset to be all data generated by OR-Instruct. Under the condition of identical training hyperparameters, we observe the performance of different Qwen2.5 model sizes on various test sets. In our study, we use the overall macro average and micro average across four test sets as metrics to evaluate the performance of the ORLMs, as visualized in Figure 10. More detailed results are provided in Appendix 8.5.2.
[Different sizes of training sets] Refer to caption [Different sizes of base models] Refer to caption
Figure 10: Scaling law validation across different data and model sizes
The figure clearly shows an upward trend in ORLM’s accuracy on the test sets as either the model size or the data volume increases. However, the impact of these two types of scaling laws on ORLM’s performance differs. Specifically, increasing the training set size results in fluctuating performance gains from the macro average metric, this is because for simple test datasets (NL4OPT and MAMO-easyLP), ORLM quickly reaches a performance bottleneck and begins to fluctuate, causing instability in its accuracy across the entire dataset. For micro average metric, the overall performance continues to exhibit a monotonic increase with the expansion of training data, approximately following a power-law trend. This indicates that increasing the amount of training data remains beneficial for improving the model’s capabilities.
Regarding the scaling law for model size, it is evident that increasing the model size significantly enhances ORLM’s accuracy. This increase not only yields steady incremental gains but also follows a power-law trend, consistent with the findings reported by Kaplan et al. (2020). This indicates that the intrinsic capability of the base model establishes a higher performance ceiling for ORLMs in optimization modeling tasks.
5.4 Inference Techniques
In previous model evaluations, we employed a greedy inference strategy to ensure the replicability and fairness of the results. However, using greedy sampling entails always selecting the word with the highest probability from the large model’s output during each generation. While this approach guarantees deterministic results, it also limits the model’s diversity and creativity. To address this, in this section, we will utilize more common sampling techniques to assess the performance improvements of the ORLMs. Specifically, we adjust the following two parameters: First, the temperature parameter controls the diversity of the generated text. Higher temperature values increase the randomness and creativity of the generated content, while lower values tend to produce more deterministic text. Second, the top-P parameter restricts the range of vocabulary choices considered by the model during generation. A higher top-P value means the model will consider a broader set of candidate words, thus enhancing the diversity of the generated content. To avoid excessive randomness that may lead to significant fluctuations in the accuracy of the generated results, we apply Pass@$k$ ($k=2,4,8$) strategy to enhance stability. Specifically, for each query, the large model generates top $k$ candidate answers, and as long as at least one of the top $k$ answers is correct, it is considered to have answered this question correctly (Pass@$k$). This approach helps improve the consistency and stability of the generated outcomes. The test results is detailed in Table 7.
Table 7: Performance of ORLM-LLaMA-3-8B on different decoding methods
Model Inference Settings NL4OPT MAMO EasyLP MAMO ComplexLP IndustryOR Micro Avg Macro Avg
ORLM-LLaMA-3-8B pass@1, temp=0, top-P=1.0 85.7% 82.3% 37.4% 38.0% 71.4% 60.8%
ORLM-LLaMA-3-8B pass@2, temp=0.7, top-P=0.95 88.6% 83.7% 49.8% 40.0% 75.6% 65.5%
ORLM-LLaMA-3-8B pass@4, temp=0.7, top-P=0.95 91.4% 85.9% 56.9% 44.0% 78.9% 69.6%
ORLM-LLaMA-3-8B pass@8, temp=0.7, top-P=0.95 93.0% 88.4% 72.1% 49.0% 83.6% 75.6%
Significant improvements have been observed in ORLMs across four test sets, with the overall micro average of Pass@8 increasing by 17.09% relative to the previous version, while the macro average rose by 24.34%. A comparison with Table 2 reveals that, in terms of both macro average and micro average accuracy, the ORLM under the Pass@8 strategy has already surpassed the accuracy of senior undergraduate students and is approaching that of experts. This highlights the significant potential of ORLMs to serve as a viable alternative to human engineers. The high Pass@8 score indicates that the correct answer is often found within the top 8 solutions. However, the model’s greedy decoding strategy remains insufficiently effective to consistently prioritize the optimal answer. The notable progress between the Pass@1 and Pass@8 scores suggests that while the model has the potential to generate correct solutions, its ranking capability is currently inadequate to reliably place the optimal solution in the first position. Fortunately, reinforcement learning may offer a promising solution to bridge the gap between Pass@1 and Pass@8, we will discuss it in the future work (See Section 6.2).
5.5 Limitations Analysis
Building upon the previous discussion and results, this section concludes with an analysis of the current limitations of ORLM. First, using ORLM-Llama-3-8B as a case study, we present a statistical breakdown of error types under the greedy inference strategy, as shown in Table 8.
Table 8: Error frequency across different datasets
Error Type NL4OPT MAMO Easy MAMO Complex IndustryOR Overall Percentage Error Percentage
Code Error 1.04% 0.00% 16.59% 31.00% 5.51% 19.77%
Model Error 12.46% 17.64% 45.97% 31.00% 22.28% 80.23%
In the table above, the "Overall Percentage" represents the occurrence frequency of each error type across all test sets, while the "Error Percentage" indicates the proportion of each error type within the total number of errors. The data reveals that, ORLM has demonstrated considerable proficiency in utilizing COPT, achieving a pass rate of approximately 95% across more than 1,000 test cases. Consequently, the primary challenges lie in the optimization modeling phase of the process.
To analyze the types of errors occurring in ORLM modeling, we randomly selected 100 problems from those where ORLM failed. Following manual evaluation and statistical analysis, and after excluding errors caused by coding issues, we categorized the modeling-related errors into three primary types, as detailed in Table 9. Notably, the predominant cause of ORLM’s modeling errors is "Low Model Completeness", which accounts for more than half of the observed errors. Further analysis reveals that this issue stems from the limitations in the expressive and learning capabilities of 8B-scale models. These models often produce overly simplified outputs in complex scenarios, reducing the intricacy of optimization modeling problems even when such scenarios are included in the training set. This highlights the importance of fine-tuning larger-scale models, a conclusion supported by our scaling law experiments.
Table 9: Analysis of main mistakes in optimization modeling
Modeling Error Type     Description Proportion
Semantic Misunderstanding Misunderstand the problem that the optimization model needs to solve. 13.40%
Errors in Objective/ Constraint Translation Correctly understand the requirements, but make errors when formulating the optimization model. 30.30%
Low Model Completeness Ignore some constraints or implicit conditions of the real problem, simplifying the complex problem. 56.30%
Although ORLM still has certain limitations, as demonstrated by previous experimental results, we recognize substantial opportunities for improvement through the application of scaling laws, different inference techniques and others. ORLM is still promising in addressing complex optimization modeling problems, even certain real-world industrial problems within specific domains. In the next section, we will demonstrate how ORLM can be applied to practical scenarios such as industrial production, education, and other real-world applications.
@@ 6 Future Direction
In previous sections, we conduct a detailed analysis of ORLM’s strengths and limitations within the field of optimization modeling. In this section, we build on above experimental results to discuss the practical applications of ORLM and potential directions for future improvement.
6.1 Potential Applications
@@
Large language models are increasingly influential across various industries, and the rise of privatized and domain-specific LLMs presents new topics for management. This work represents the first attempt to fine-tune open-source LLMs specifically for optimization modeling. The ORLM not only surpasses the performance of cutting-edge proprietary models but also ensures privacy, security, and stability for commercial applications. Additionally, its 7B size allows for private deployment on personal computers with only 16GB of VRAM, making operations research more accessible to a wider audience. We hope that this work could shed lights on the deep integration of large language models with operations research and management science. Potential applications are listed as follows:
  1. 1.
    Revolutionizing the role of operations research in industry: In recent years, operations research has rapidly permeated various industrial sectors, supporting companies in making more informed and effective decisions. However, optimization modeling tasks in industry often encounter two critical challenges: (1) optimization models are typically not robust enough to withstand changes in external conditions. When the environment shifts, models must be quickly adjusted to help companies make timely decisions under new circumstances; (2) optimization projects typically rely on business experts to specify goals and constraints, while algorithm engineers are responsible for developing and solving the models. For many specialized industrial problems, communication between these two roles can be costly. Algorithm engineers may struggle to grasp the intricacies of business logic, and business experts often lack technical modeling expertise.
    The introduction of ORLM could help address many of these challenges. Particularly, companies with extensive training data can customize ORLMs to their domain based on OR-Instruct. As demonstrated in Tables 2 and 3, ORLM significantly enhances optimization modeling capabilities while retaining the broad knowledge acquired during pre-training. This allows it to quickly grasp industry-specific terminology and facilitate effective communication between business experts and algorithm engineers. ORLM operates with remarkable efficiency, generating an initial solution within seconds that both engineers and business teams can refer to. Given ORLM’s strong performance on Pass@8 metrics, it can also produce multiple solutions, supporting selection and comparison, which accelerates project iterations. Two representative workflows are illustrated in Figure 11.
    [In the course of enterprise development, the constraints and objectives of operational contexts continually evolve, necessitating swift adjustments to existing algorithmic models to meet business demands.] Refer to caption [For newly arising decision-making problems, enterprises can leverage ORLM for rapid modeling and experiment with small data samples. This approach enables enterprises to respond swiftly to market changes.] Refer to caption
    Figure 11: Utilize ORLM for adaptive modeling in enterprise decision-making.
    Moreover, ORLM can seamlessly modify existing optimization models by adding new business constraints, creating preliminary models, and generating initial solution code, which engineers can then validate. This capability not only reduces delivery costs but also improves accuracy on complex tasks.
    Beyond these capabilities, ORLM has the potential to leverage Retrieval-Augmented Generation (RAG) and similar technologies to incorporate internal company documents, evolving into a domain-specific optimization expert well-versed in the company’s field. This feature allows ORLM to support knowledge transfer and updates, transforming traditional project reports into valuable data assets for future model training. In addition to project delivery, ORLM can also be applied in training junior algorithm engineers in modeling techniques, further enhancing its value within the enterprise.
  2. 2.
    Operations research education: The outstanding performance of ORLM in fundamental optimization modeling problems makes it sufficiently capable of addressing the modeling challenges encountered by beginners in operations research. Instructors can utilize ORLM for efficient instructional guidance, allowing students to quickly derive reference solutions and processes for unfamiliar operations research modeling problems, thereby enhancing teaching efficiency and quality.
  3. 3.
    Mathematical modeling competitions: As one of the most influential competitions, mathematical modeling contests aim to cultivate students’ abilities to solve problems using mathematical knowledge. Since 2023, several mathematical modeling competitions have begun to acknowledge the significant role of generative AI in modeling, permitting its use as an aid (e.g., MCM 2024). The capabilities of ORLM position it as a powerful tool in such competitions.
  4. 4.
    Reducing the learning cost of solvers: Solvers are a core component of optimization modeling problems, yet the diverse programming syntax of different solvers significantly increases the learning cost for users, hindering their widespread adoption. The emergence of ORLM marks a pivotal shift in this regard. Notably, Table 8 indicates that ORLM rarely encounters errors due to coding issues (95% accuracy rate), even with less common syntax such as that specific to COPT. This suggests that algorithm engineers can confidently develop reliable operations research models and rely on ORLM to translate them into corresponding solver code, alleviating the need to extensively learn solver syntax.
6.2 Future Research Directions
In Section 5, we conduct extensive experiments to analyze the current limitations of ORLM and to investigate the underlying causes. Based on our findings, we propose directions for future improvements, calling for community attention and collaboration to drive progress in this field:
  • Incorporating sophisticated techniques: In Section 5.4, we observe that while ORLM can generate correct answers, it struggles with ranking them effectively. Incorporating reinforcement learning techniques could help narrow the gap between pass@1 and pass@8 scores. As noted by Sutton and Barto (2018), one of the primary goals of reinforcement learning is to align the policy with the Bellman optimality policy, effectively optimizing ranking quality. In this context, relevant methods like reinforcement learning from human feedback (RLHF), direct preference optimization (DPO), and Kahneman-Tversky optimization (KTO) aim to minimize the disparity between the model’s current greedy decoding policy and the Bellman optimality policy, thereby enhancing the model’s ranking accuracy and moving pass@1 closer to pass@8. In addition, approaches such as multi-agent collaboration and internal chain of thought structures similar to o1 (OpenAI 2024) could further enhance the performance of current large models in optimization modeling tasks.
  • Dataset construction: While this paper presents a data synthesis strategy, it is insufficient for reinforcement learning techniques like RLHF, which require training data in the form of a preference list. For each question, multiple responses must be ranked according to preference. Furthermore, in the context of optimization problems, incorporating actual optimal solutions into the training set would enable a wider range of techniques. Constructing such datasets may require user’s feedback or the development of new data synthesis methods to better meet these needs.
  • Data mining and exploration: As shown in Figure 10, we have demonstrated that scaling laws significantly benefit ORLM. However, for a fixed model size, the performance gains from additional data become increasingly marginal at later stages. This raises the question of whether it is possible to identify an optimized, minimal dataset or develop efficient data synthesis strategies that can balance ORLM’s performance with the associated training costs. Existing studies (e.g., Chen et al. 2023b) indicate that only a small portion of a complete dataset is often sufficient to achieve comparable performance. Therefore, developing a tailored data exploration strategy for optimization problem modeling, which minimizes the reliance on large-scale data resources for enhancing large models, represents an important direction for future research.
@@ 7 Conclusion
In addition, this paper reveals the current limitations of ORLM, such as its overly simplistic output for complex problems, weak ability to rank optimal solutions, and poor learning capability from data. Our experimental results suggest that ORLM can be further enhanced by following scaling laws. For example, fine-tuning larger models with superior overall capabilities can enhance ORLM’s performance. Additionally, increasing the quantity and quality of training data or adjusting the data composition could bolster ORLM’s ability to address certain weaker problems. Furthermore, methods such as Reinforcement Learning from Human Feedback (RLHF), which train models based on preference data, could be introduced to further improve ORLM’s ranking ability for optimal solutions, potentially reaching expert-level performance. This finding offers significant insights for advancing ORLM’s applications in industry and highlights promising directions for future research.